Can you please explain fraction equations to me? Example- 1/2(a+4)=24. thank you

Answer by njusa(12)

The first step in a equation such as this is to remove the fraction. In order to do this we have to multiply both sides of the equation by 2.
2*1/2(a+4)=24*2
1(a+4) = 48
a+4 = 48
a = 44